# The Degree of (Extended) Justified Representation and Its Optimization

### v1

@@ 2.1 Proof of Theorem 1 and 2
From the proof of NP-hardness, we further have MDEJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}}{\bar{n}+\bar{m}}$ (Note that $\bar{n}$ and $\bar{m}$ are the parameters of the sparse-SAT problem rather than ABC voting, we will reformulate it later). Now, instead of reducing from sparse-SAT problems directly, we add an intermediate reduction. Given a sparse-SAT instance $\phi$ with $\bar{n}$ clauses and $\bar{m}$ different variables (say $x_{1},\dots,x_{\bar{m}}$), there is a corresponding ‘overstaffing’ sparse-SAT instance $\phi^{\prime}$ with $\bar{n}^{\prime}=\bar{n}+1$ clauses and $\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ different variables such that a new clause is added with a number of $(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ new variables like $(y_{1}\vee y_{2}\vee\dots)$. Obviously, the ‘overstaffing’ sparse-SAT problem is also an NP-hard problem.
@@
Now we reduce our problem from ‘overstaffing’ sparse-SAT problem in the same way that we reduce from sparse-SAT problem. Similarly, MDEJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}^{\prime}}{\bar{n}^{\prime}+\bar{m}^{\prime}}$ where $\bar{n}^{\prime}=\bar{n}+1,\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$. Recall the construction, the number of voters and the committee size are $n_{\text{voting}}=\bar{n}^{\prime}\bar{m}^{\prime}+\bar{m}^{\prime 2}+\bar{n}^{\prime}+\bar{m}^{\prime}$ and $k_{\text{voting}}=\bar{m}^{\prime}+1$, respectively, which can be reformulated as $\frac{k_{\text{voting}}}{n_{\text{voting}}}=\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}$. Since
$(\bar{n}^{\prime}+\bar{m}^{\prime})^{\epsilon}=(\bar{n}^{\prime}+\bar{m}+(\bar{n}^{\prime}+\bar{m})^{\lceil\frac{1}{\epsilon}\rceil})^{\epsilon}>\bar{n}^{\prime}+\bar{m}>\bar{n}^{\prime}.$
Finally, MDJR cannot be approximated in polynomial time to within a factor of
@@
where $\epsilon>0$. ∎
@@ 3.2.1 Algorithm for MDJR
Our starting point is the algorithm GreedyAV (Algorithm 1) proposed by Aziz et al. [2], which always outputs a JR committee.
Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
Output: A winning committee $W$ of size $k$
1 $W\leftarrow\emptyset$
2for $j\in[k]$ do
3       Let $c$ be the candidate approved by the maximum number of voters in $N$
4      $W\leftarrow W\cup c$
5      $N\leftarrow N\setminus V$ where $V$ is the set of voters who approve $c$
return $W$
Algorithm 1 Greedy Approval Voting (GreedyAV)
Next, we show the following property for GreedyAV which is the key for our algorithm. It states that the algorithm GreedyAV also gives us the optimal JR degree if $n$ is large enough.
@@ 5      $N\leftarrow N\setminus V$ where $V$ is the set of voters who approve $c$
Given any instance with the maximum achievable JR degree of $c_{\emph{max}}$ and $n>k^{2}(c_{\emph{max}}-1)$, GreedyAV will output the committee achieving JR degree $c_{\emph{max}}$.
@@
We prove it by contradiction. Suppose the committee output by GreedyAV provides the JR degree less than $c_{\text{max}}$. Hence, there will be at least $(\frac{n}{k}-c_{\text{max}}+1)$ voters that approve a non-elected candidate. From the definition of GreedyAV, the coverage of voters is at least $(\frac{n}{k}-c_{\text{max}}+1)$ in each iteration, and at least $\frac{n}{k}$ in the first iteration. Therefore, the total number of voters is at least
$\frac{n}{k}+(k-1)\left(\frac{n}{k}-c_{\text{max}}+1\right)+\left(\frac{n}{k}-c_{\text{max}}+1\right)$
$=n+\frac{n}{k}-k(c_{\text{max}}-1)>n,$
which leads to a contradiction in the number of voters. ∎
Given any instance with $n>k^{2}(c_{\text{max}}-1)$, GreedyAV can achieve the maximum JR degree. Hence, the remaining case is $n\leq k^{2}(c_{\text{max}}-1)$. Given any instance $(N,C,\mathbf{A},k)$, each candidate $c_{i}\in C$ can be seen as a subset of $N$, including all voters that approve $c_{i}$. Hence, there are at most $2^{n}$ different types of candidates, implying that every instance corresponds to an equivalent instance with $m\leq 2^{n}$. Therefore, we can decide whether there exists a committee providing JR degree of $c$ by enumerating all the committees when $n\leq k^{2}(c-1)$, which can be computed with running time
@@ 3.2.2 Algorithm for MDEJR
$s_{\text{PAV}}(W)=\sum_{i=1}^{n}\sum_{j=1}^{|A_{i}\cap W|}\frac{1}{j}.$
Input: An instance $\mathcal{I}=(N,C,\mathbf{A},k))$
Output: A winning committee $W$ of size $k$
1 $W\leftarrow k$ arbitrary candidates from $C$
3       $W\leftarrow W\setminus\{c^{-}\}\cup\{c^{+}\}$
return $W$
Algorithm 3 $\lambda$-LS-PAV
Next, we prove the following observation in a similar spirit to Proposition 2, which shows that the local search variant of PAV can achieve maximum EJR degree if $n$ is sufficiently large and $\lambda$ is sufficiently small.
@@ 3       $W\leftarrow W\setminus\{c^{-}\}\cup\{c^{+}\}$
Given any instance with the maximum achievable EJR degree of $c_{\emph{max}}$. For all $\lambda<\frac{n}{k^{2}}$ satisfying $n\geq k(k+1)(c_{\emph{max}}-1)+\lambda k^{2}$, $\lambda$-LS-PAV will output the committee achieving EJR degree $c_{\emph{max}}$.
@@
We prove it by contradiction. Suppose the committee, $W$, output by $\lambda$-LS-PAV provides the EJR degree strictly smaller than $c_{\text{max}}$. There exists a $\ell$-cohesive group $V\subseteq N$, such that less than $c_{\text{max}}$ voters in $V$ approves at least $\ell$ members of $W$, i.e., |$|\{i\in V:|A_{i}\cap W|\geq\ell\}|<c_{\text{max}}$. Since $V$ is $\ell$-cohesive, there exist $\ell$ candidates approved by all voters in $V$. At least one such candidate, $c^{+}\in\bigcap_{i\in V}A_{i}$, is not selected, as otherwise all voters in $V$ approve at least $\ell$ members of $W$.
We will show that there exists a candidate $c^{-}\in W$ that the increment of the score by swapping $c^{+}$ and $c^{-}$ is at least $\lambda$, so $W$ cannot be an output of $\lambda$-LS-PAV (since the $while$ loop can be executed continually), a contradiction. To see this, we try to swap $c^{+}$ and any candidate $c^{-}\in W$. Since $c^{+}\in A_{i}$ for all $i\in V$, we have
$\Delta(W,c^{+},c^{-})$
$=$ $\sum_{\begin{subarray}{c}i:c^{+}\in A_{i}\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{\begin{subarray}{c}i:c^{+}\notin A_{i}\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
and, by summing up $\Delta(W,c^{+},c^{-})$ for $c^{-}\in W$,
$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})$
$\geq$ $\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in V:\\ c^{-}\notin A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|+1}-\sum_{c^{-}\in W}\sum_{\begin{subarray}{c}i\in N\setminus V:\\ c^{-}\in A_{i}\end{subarray}}\frac{1}{|A_{i}\cap W|}$
$=$ $\sum_{i\in V}\frac{k-|A_{i}\cap W|}{|A_{i}\cap W|+1}-n+|V|=\sum_{i\in V}\frac{k+1}{|A_{i}\cap W|+1}-n$
$\geq$ $\!\!\!\!\!\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{|A_{i}\cap W|+1}-n\geq\!\!\!\!\!\sum_{i\in V:|A_{i}\cap W|<\ell}\!\frac{k+1}{\ell}-n.$
Since less than $c_{\text{max}}$ voters in $V$ approves at least $\ell$ members of $W$, the number of voters in $V$ approves less than $\ell$ members of $W$ is at least $|V|-c_{\text{max}}+1$. Thus,
$\sum_{c^{-}\in W}\Delta(W,c^{+},c^{-})$ $\geq\sum_{i\in V:|A_{i}\cap W|<\ell}\frac{k+1}{\ell}-n$
$\geq(k+1)\frac{|V|-c_{\text{max}}+1}{\ell}-n$
$\geq(k+1)(\frac{n}{k}-c_{\text{max}}+1)-n$
$=\frac{n}{k}-(k+1)(c_{\text{max}}-1)\geq\lambda k.$
From the pigeonhole principle it follows that there exists a candidate $c^{-}\in W$ such that $\Delta(W,c^{+},c^{-})\geq\lambda$. ∎
Since Proposition 3 holds for every initial committee $W$ in $\lambda$-LS-PAV. By considering the initial committee $W$ being the one with the maximum PAV-score and $\lambda=0$ (in which case the while-loop is never executed), we have the following corollary. In contrast to our counterexample in Appendix C, this corollary shows the success of PAV for large enough $n$, which may be of independent interest.
@@
Given any instance with the maximum achievable EJR degree of $c_{\emph{max}}$. If $n\geq k(k+1)(c_{\emph{max}}-1)$, PAV has an EJR degree of $c_{\emph{max}}$.
@@
Given any instance $(N,C,\mathbf{A},k)$ with maximum achievable EJR degree $c_{\emph{max}}$. If $n\geq k(k+1)^{2}(c_{\emph{max}}-1)$, $\frac{n}{k(k+1)}$-LS-PAV outputs a committee with EJR degree $c_{\emph{max}}$.
Based on Corollary 2, we can use a similar way to MDJR to design our algorithm. In particular, we can decide whether there exists a committee providing EJR degree $c$ by enumerating all the committees when $n\leq k(k+1)(c_{\text{max}}-1)$, which can be computed in running time $f(k,c)$. Our algorithm is presented in Algorithm 4, which achieves the maximum EJR degree and runs in time $f(k,c_{\text{max}})\cdot\text{poly}(m,n)$. The correctness of our algorithm follows from arguments similar to MDJR. The time complexity analysis is also similar, except that we need to show $\frac{n}{k(k+1)}$-LS-PAV runs in polynomial time. To see this, each swap operation can be executed in polynomial time with a minimum score increment by $\lambda=\frac{n}{k(k+1)}$, and the maximum PAV-score is $n\cdot(1+\frac{1}{2}+\cdots+\frac{1}{k})=O(n\ln k)$. Thus, the number of while-loop executions is bounded by $O(k^{2}\ln k)$.
@@ 1 $W\leftarrow\frac{n}{k(k+1)}\text{-LS-PAV}(\mathcal{I})$
2for $c:\lceil\frac{n}{k(k+1)}\rceil$ to $\lfloor\frac{n}{k}\rfloor$ do
@@ 4 Conclusion and Future Work
Many potential further works can be explored. For example, one can explore whether the negative results can be circumvented by considering restricted domains of preferences [13, 37]. Another direction is to consider some other quantitative measurements with respect to (E)JR, e.g., using the ratio instead of the number. As we have discussed earlier, our definition of EJR degree using numbers is more aligned with the original definition of EJR, whereas the definition using ratios/fractions, by prioritizing more on $\ell$-cohesive groups with large $\ell$, is more aligned with the notion of Individual Representation [7] discussed in Sect. 1.3. We believe both the “number version” and the “ratio version” are worth-studied. Which choice is better depends on the specific applications.

### v2

@@ 2.2 Matching Inapproximability Results
From the proof of NP-hardness, we have that MDJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}}{\bar{n}+\bar{m}}$, where $\bar{n}$ is the number of the clauses and $\bar{m}$ is the number of the variables. Now, instead of reducing from sparse-SAT problems directly, we add an intermediate reduction. Given a sparse-SAT instance $\phi$ with $\bar{n}$ clauses and $\bar{m}$ different variables, we construct another sparse-SAT instance $\phi^{\prime}$ with $\bar{n}^{\prime}=\bar{n}+1$ clauses and $\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ different variables such that a new clause is added with $(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$ new variables. Obviously, $\phi$ and $\phi^{\prime}$ have the same satisfiability.
@@
Now, we use $\phi^{\prime}$ instead of $\phi$ to construct the ABC voting instance. By our previous analysis, MDJR cannot be approximated in polynomial time to within a factor of $\frac{\bar{n}^{\prime}}{\bar{n}^{\prime}+\bar{m}^{\prime}}$ where $\bar{n}^{\prime}=\bar{n}+1,\bar{m}^{\prime}=\bar{m}+(\bar{n}+\bar{m}+1)^{\lceil\frac{1}{\epsilon}\rceil}$. Recall that the number of voters and the committee size are $n_{\text{voting}}=\bar{n}^{\prime}\bar{m}^{\prime}+\bar{m}^{\prime 2}+\bar{n}^{\prime}+\bar{m}^{\prime}$ and $k_{\text{voting}}=\bar{m}^{\prime}+1$, respectively, which can be reformulated as $\frac{k_{\text{voting}}}{n_{\text{voting}}}=\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}$. Since
$(\bar{n}^{\prime}+\bar{m}^{\prime})^{\epsilon}=\left(\bar{n}^{\prime}+\bar{m}+(\bar{n}^{\prime}+\bar{m})^{\lceil\frac{1}{\epsilon}\rceil}\right)^{\epsilon}>\bar{n}^{\prime}+\bar{m}>\bar{n}^{\prime},$
MDJR cannot be approximated in polynomial time to within a factor of
@@
where $\epsilon>0$.
As we have mentioned, the same analysis holds for MDEJR, as there is no $\ell$-cohesive group with $\ell\geq 2$ in our construction. ∎
Proof of Theorem 5 and 6.
Both theorems can be proved with only minor modifications to the last step in the proof of Theorem 3 and 4 above:
$\left(\frac{1}{\bar{n}^{\prime}+\bar{m}^{\prime}}\right)^{1-\epsilon}<\left(\frac{1}{1+\bar{m}^{\prime}}\right)^{1-\epsilon}=\left(\frac{1}{k_{\text{voting}}}\right)^{1-\epsilon}.\qed$
@@ 3.2.1 Algorithm for MDJR
Our starting point is the algorithm GreedyAV (Algorithm 1) proposed by Aziz et al. [2]. We show the following property for GreedyAV which is the key for our algorithm. It states that the algorithm GreedyAV also gives us the optimal JR degree if $n$ is large enough.
@@
Given any instance with the maximum achievable JR degree of $c_{\max}$ and $n>k^{2}(c_{\max}-1)$, GreedyAV will output a committee achieving JR degree $c_{\max}$.
@@
$n>k^{2}(c_{\max}-1)$ implies $c_{\max}<\frac{n}{k^{2}}+1$. This proposition then follows from Proposition 2. ∎
Given any instance with $n>k^{2}(c_{\text{max}}-1)$, GreedyAV can achieve the maximum JR degree. Hence, the remaining case is $n\leq k^{2}(c_{\text{max}}-1)$. Given any instance $(N,C,\mathbf{A},k)$, each candidate $c_{i}\in C$ can be seen as a subset of $N$ including all voters that approve $c_{i}$. Hence, there are at most $2^{n}$ different types of candidates, implying that every instance corresponds to an equivalent instance with $m\leq 2^{n}$. Therefore, we can decide whether there exists a committee providing JR degree of $c$ by enumerating all the committees when $n\leq k^{2}(c-1)$, which can be computed with running time
@@ 3.2.2 Algorithm for MDEJR
For MDEJR, we prove the following observation in a similar spirit to Proposition 4, which shows that the local search variant of PAV (Algorithm 2) can achieve maximum EJR degree if $n$ is sufficiently large and $\lambda$ is sufficiently small.
@@
Given any instance with the maximum achievable EJR degree of $c_{\max}$. For any $\lambda\in[0,\frac{n}{k^{2}})$ satisfying $n>k(k+1)(c_{\max}-1)+\lambda k^{2}$, $\lambda$-LS-PAV will output the committee achieving EJR degree $c_{\max}$.
@@
$n>k(k+1)(c_{\max}-1)+\lambda k^{2}$ implies $c_{\max}<\frac{n}{k(k+1)}-\lambda\frac{k}{k+1}+1$. Proposition 3 implies this proposition. ∎
Since Proposition 5 holds for every initial committee $W$ in $\lambda$-LS-PAV, by considering the initial committee $W$ being the one with the maximum PAV-score and $\lambda=0$ (in which case the while-loop is never executed), we have the following corollary. In contrast to our counterexample in Appendix C, this corollary shows the success of PAV for large enough $n$, which may be of independent interest.
@@
Given any instance with the maximum achievable EJR degree of $c_{\max}$. If $n>k(k+1)(c_{\max}-1)$, PAV has an EJR degree of $c_{\max}$.
@@
Given any instance $(N,C,\mathbf{A},k)$ with maximum achievable EJR degree $c_{\max}$. If $n>k(k+1)^{2}(c_{\max}-1)$, $\frac{n}{k(k+1)}$-LS-PAV outputs a committee with EJR degree $c_{\max}$.
Based on Corollary 2, we can use a similar way to MDJR to design our algorithm. In particular, we can decide whether there exists a committee providing EJR degree $c$ by enumerating all the committees when $n\leq k(k+1)^{2}(c_{\max}-1)$, which can be computed in running time $f(k,c)$. Our algorithm is presented in Algorithm 4, which achieves the maximum EJR degree and runs in time $f(k,c_{\max})\cdot\text{poly}(m,n)$. The correctness of our algorithm follows from arguments similar to MDJR. The time complexity analysis is also similar, except that we need to show $\frac{n}{k(k+1)}$-LS-PAV runs in polynomial time. To see this, each swap operation can be executed in polynomial time with a minimum score increment by $\lambda=\frac{n}{k(k+1)}$, and the maximum PAV-score is $n\cdot(1+\frac{1}{2}+\cdots+\frac{1}{k})=O(n\ln k)$. Thus, the number of while-loop executions is bounded by $O(k^{2}\ln k)$.
@@ 1 $W\leftarrow\frac{n}{k(k+1)}\text{-LS-PAV}(\mathcal{I})$
2for $c:\lceil\frac{n}{k(k+1)^{2}}\rceil$ to $\lfloor\frac{n}{k}\rfloor$ do
@@ 4 Conclusion and Future Work
Many potential further works can be explored. For example, one can explore whether the negative results can be circumvented by considering restricted domains of preferences [13, 37]. Another direction is to consider some other quantitative measurements with respect to (E)JR, e.g., using the ratio instead of the number. As we have discussed earlier, our definition of EJR degree using numbers is more aligned with the original definition of EJR, whereas the definition using ratios/fractions, by prioritizing more on $\ell$-cohesive groups with large $\ell$, is more aligned with the notion of Individual Representation [7] discussed in Sect. 1.3. We believe both the “number version” and the “ratio version” are worth studying. Which choice is better depends on the specific applications.
Acknowledgments
The research of Biaoshuai Tao was supported by the National Natural Science Foundation of China (No. 62472271) and the Key Laboratory of Interdisciplinary Research of Computation and Economics (Shanghai University of Finance and Economics), Ministry of Education.
The authors gratefully appreciate the valuable comments from IJCAI’24, AAAI’25, and AAMAS’25 reviewers, which are very helpful in improving this paper. We especially appreciate that one of the AAMAS’25 reviewers points out the $\frac{1}{k}$-approximation of the greedy algorithm, which leads to Theorem 1 and 2 in this paper.